x - square root of x=0
Answer by AnlytcPhil(1806)

x - square root of x = 0 
         _
    x - Öx = 0

Isolate the radical term
              _
         x = Öx

Square both sides:
               _
      (x)2 = (Öx)2

        x2 = x

 Get 0 on right by subtracting x from both sides

    x2 - x = 0

Factor left side

  x(x - 1) = 0

Set each factor = 0

x = 0 is its own solution

x - 1 = 0 which has solution x = 1

Now we must always check the answers to radical problems to 
make sure they are solutions.
Checking x = 0
         _
    x - Öx = 0
         _
    0 - Ö0 = 0
    
        0 = 0 

Checking x = 1
         _
    x - Öx = 0
         _
    1 - Ö1 = 0
    
        0 = 0

Edwin
AnlytcPhil@aol.com